The activation energy of a reaction is 22.5 k Cal mol-1 and the value of rate constant at 40°C is 1.8 ×10−5 s−1. Calculate the frequency factor, A.


1
Expert's answer
2021-04-01T02:34:59-0400

"Ea = 22.5 kcal mol^{-1}= 22500 cal mol^{-1}"


"T= 40\u00b0C = 40 + 273 = 313 K"


"k = 1.8 \u00d710^{-5}sec ^{-1}"


Substituting the values in the equation


"log A = log k + (\frac{E_a}{2.303RT})"


"log A = log (l.8×10^{-5}) + (\frac{E_a}{2.303×1.987×313})"


"log A = log (l.8) \u2013 5 + (15.7089)"


"log A = (10.9642)"


"A = antilog ( 10.9642)"


"A = 9.208 \u00d7 10^{10}" collisions "s^{-1}"